The Butler-Perkins Company (BPC) must decide between two mutually exclusive projects. Each costs $6,500 and has an expected lB.) Based on the risk-adjusted NPVs, which project should BPC choose? Project A or B? C.) If you knew that Project B's cash flows were negatively correlated with the firm's other cash flow, but Project A's cash flows were positively correlated, how might this affect the decision? Would it make Project B more or less appealing? D.) If Project B's cash flows were negatively correlated with gross domestic product (GDP), while A's cash flows were positively correlated, would that influence your risk assessment? Would it make project B more or less appealing?
Answer #1

a. Project A's expected annual cash flow= 0.2*6000+0.6*6500+0.2*7000=$6500

Project B's expected annual cash flow= 0.2*0+0.6*6500+0.2*19000=$7700

Standard deviation of A=sqrt of (0.2*(6000-6500)^2+0.6*(6500-6500)+0.2*(7000-6500)^2)=$316.22

Coefficient of variation of A= 316.22/6500=0.0486

B. As the Coefficient of variation is lower for A, BPC should choose Project A

This will not change, the decision and Project A will still be more attractive than B.

Relationship with GDP also will not change the decision as we have considered all such probability in the cash flow determination.